# Two Types of AI Existential Risk: Decisive and Accumulative

### v2

@@ Abstract
The conventional discourse on existential risks (x-risks) from AI typically focuses on abrupt, dire events caused by advanced AI systems, particularly those that might achieve or surpass human-level intelligence. These events have severe consequences that either lead to human extinction or irreversibly cripple human civilization to a point beyond recovery. This discourse, however, often neglects the serious possibility of AI x-risks manifesting incrementally through a series of smaller yet interconnected disruptions, gradually crossing critical thresholds over time. This paper contrasts the conventional decisive AI x-risk hypothesis with an accumulative AI x-risk hypothesis. While the former envisions an overt AI takeover pathway, characterized by scenarios like uncontrollable superintelligence, the latter suggests a different causal pathway to existential catastrophes. This involves a gradual accumulation of critical AI-induced threats such as severe vulnerabilities and systemic erosion of econopolitical structures. The accumulative hypothesis suggests a boiling frog scenario where incremental AI risks slowly converge, undermining resilience until a triggering event results in irreversible collapse. Through systems analysis, this paper examines the distinct assumptions differentiating these two hypotheses. It is then argued that the accumulative view reconciles seemingly incompatible perspectives on AI risks. The implications of differentiating between these causal pathways — the decisive and the accumulative — for the governance of AI risks as well as long-term AI safety are discussed.
@@ 1 Introduction
Recent advances in machine learning have sparked intense debate about the existential risks (x-risks) associated with Artificial intelligence (AI) systems. Central to this debate is a concern about the potential pathways through which AI could cause existential catastrophes. In direct response to this concern, this paper aims to explore: How should we effectively conceptualize AI x-risks in the context of various types of causal pathways leading to AI-induced existential catastrophes? Conventional discourse on AI existential catastrophes typically portrays them as sudden, decisive events, often triggered by artificial general intelligence or artificial superintelligence (e.g., Bostrom (2013) and Ord (2020a)).
Contrasting this traditional decisive viewpoint, this paper introduces the accumulative AI x-risk hypothesis as an alternative lens. The accumulative hypothesis posits that AI x-risks do not exclusively materialize as high-magnitude events initiated by artificial general intelligence or artificial superintelligence. Instead, they can accumulate incrementally through a series of smaller yet consequential AI-induced disruptions over prolonged periods. These disruptions interact synergistically, gradually undermining systemic resilience, ultimately leading to a critical point where a stressor event could result in an irreversible collapse.
This paper defends and develops this accumulative perspective which is not adequately represented or robustly defended in the philosophical literature. I argue that a thorough and nuanced understanding of AI x-risks necessitates the integration of this accumulative hypothesis into both qualitative and quantitative x-risk studies. Recognizing the accumulative nature of AI x-risks is critical for the governance of AI risks and the development of long-term AI safety initiatives. This paper hence suggests a reevaluation of existing approaches to AI x-risks in light of the accumulative hypothesis and encourages a more diverse perspective to AI x-risk management.
@@ 2 AI x-risk: preliminaries
Existential risks are risks associated with existential catastrophic events. These events result in either the extinction of humanity or an unrecoverable decline in humanity’s potential to thrive (Ord, 2020a). [^1] Some definitions, such as the one proposed by Bostrom (2013, p. 15), broaden the scope of existential threats to include not only human life but all sentient beings: “An existential risk is one that threatens the premature extinction of Earth-originating intelligent life or the permanent and drastic destruction of its potential for desirable future development.” Despite this broader definition, this paper deliberately employs the term ”humanity” as the focal subject of AI-induced existential catastrophes (Ord, 2020a). The choice is pragmatic and is motivated by the predominant narrative in AI x-risk scholarly literature and public discourse, which has classically concentrated on the catastrophic impacts on humanity and human civilization. Additionally, this choice aligns with leading AI research laboratories such as Google DeepMind or OpenAI regarding the development of AGI or ASI for all of humanity. This terminological choice hence does not undermine the significance of non-human sentient beings. Existential catastrophes are a class of potential events that may originate from natural causes, such as a supervolcanic eruption; anthropogenic sources, like nuclear conflict; or emerging threats, such as misaligned artificial superintelligence. [^2] For a historical analysis of existential risks, see Torres (2023). This paper concentrates on existential catastrophes induced by AI (AI x-catastrophes) and the associated AI x-risks.
@@
The conventional discourse on AI x-catastrophes portrays them as decisive, large-scale events caused by highly advanced AI systems, often referred to as artificial general intelligence (AGI) or artificial superintelligence (ASI). [^3] In AI lexicon, AGI roughly represents a capability level on par with human intelligence, while ASI denotes capabilities surpassing human intelligence in achieving objectives. Typically, it is argued that AI x-risks are caused by AGI or ASI (i.e., Bostrom (2014); Ord (2020a, b)). Another perspective suggests that AI x-catastrophes might arise from AI systems lacking complete general or super intelligence, yet possessing specific extremely powerful capabilities (e.g., Carlsmith (2022)). The preference to tie x-risks to AGI or ASI, rather than extremely capable but narrow AI systems, in this paper is motivated by the following consideration: the x-risks from AGI or ASI have been employed by influential voices in the academic field of x-risk studies as well as pioneering AI research labs such as OpenAI. The decisive nature of x-risks according to the conventional discussions equally holds if we replace the term AGI or ASI with extremely dangerous capable yet narrow AI that are not generally or superly intelligent, but have very narrow-focused uncontrollable capabilities. The terminological choice for AGI or ASI x-risks, therefore, does not affect the characterization of the conventional view on AI x-risk in this paper. The idea that advanced machines can pose significant risks is not fully novel and has historical antecedents. Samuel Butler (1863, p.185), a novelist and literary critic, alluded to the possibility of machines dominating humanity. This concern was later picked up by the renowned mathematician Alan Turing (1950), who warned of intelligent machines eventually taking control. Norbert Wiener (1960), a founder of the field of cybernetics, cautioned against entrusting machines with purposes misaligned with human intentions or desires. Similarly, the mathematician Irving J. Good (1966) expressed concerns about the creation of ultraintelligent machines.
@@
More recently, philosophers like Nick Bostrom (2002, p.7) drew systematic attention to the existential threats posed by ASI: “When we create the first superintelligent entity, we might make a mistake and give it goals that lead it to annihilate humankind, assuming its enormous intellectual advantage gives it the power to do so.” [^4] While Bostrom’s work has systematized the discussion of x-risks from ASI, the original post-2000 discussions trace back to Eliezer Yudkowsky’s views concerning AGI, ASI, and their associated x-risks. Ben Goertzel (2015) explores the early conceptual evolution of these topics, tracing their roots to Yudkowsky’s initial informal explorations. I primarily reference Bostrom as a representative figure who brought systematic and philosophical depth to concepts that Yudkowsky and others initially introduced and examined in speculative media and blog posts. Computer scientists such as Stuart Russell (2019) and physicists like Max Tegmark (2018) echoed similar concerns, stressing the x-risks of ASI beyond human control. Such views about ASI x-risks frequently hinge on two key theses: orthogonality and instrumental convergence (Bostrom, 2012, 2014).
The orthogonality thesis posits an AI system’s intelligence and its goals are independent variables. This implies that an ASI could be designed with any number of objectives, benevolent or otherwise, irrespective of its intellectual capacity. That is, an AI system can be superintelligent but possess x-catastrophic goals as readily as it could have beneficial ones. The instrumental convergence thesis proposes that ASI would seek self-preservation and goal achievement by any means necessary. This could include removing obstacles like humans. Two hypothetical scenarios illustrate conventional models of causal pathway derived from the dual theses, which potentially lead to x-catastrophic outcomes from ASI.
In a thought experiment by Bostrom (2003), an ASI is given the goal of maximizing paper clip production. This AI could quickly realize that human existence is counterproductive to its goal. Humans pose a threat as they might deactivate the AI, resulting in fewer paper clips. Additionally, the AI might recognize that human bodies contain atoms that could be repurposed into paper clips. The ASI’s optimal future then involves a world abundant with paper clips but devoid of humans. In a second thought experiment, Russell and Norvig (2010, p.1039) reference a hypothetical scenario attributed to Marvin Minsky, a pioneer in AI research, where an advanced AI, tasked with solving the Riemann hypothesis, could decide to commandeer Earth’s resources to construct supercomputers, thereby aiding its objective and endangering humanity’s existence. The two hypothetical examples illustrate how ASI could independently pursue its goals to the detriment of humanity, and could deliberately or accidentally cause x-catastrophes, particularly when humans are unable to effectively steer the superintelligence’s behavior. Both experiments have in common the following causal pathway: a superintelligent reward-maximizer paired with a goal system that gives rewards for achieving existentially-catastrophic goals.
@@
On the conventional view sketched above, the path to the occurrence of AI x-catastrophes is via conclusive actions of ASI. Toby Ord (2020a, p. 20), a notable figure in x-risk studies, explicitly endorses the decisive character of x-catastrophes: “I take on the usual sense of catastrophe as a single decisive event rather than any combination of events that is bad in sum. A true existential catastrophe must by its very nature be the decisive moment of human history the point where we failed.” Given that AI x-risks concern the potential for AI x-catastrophes, the conventional viewpoint suggests that AI x-risks are about sudden, cataclysmic events that could either eradicate humanity or significantly curtail its potential. The core assumption underlying this viewpoint can be articulated as the decisive ASI x-risk hypothesis:
@@
Bostrom’s portrayal of ASI incorporating extremely dangerous goals leading to the annihilation of humanity points at a conclusive event with extremely dire consequences. Similarly, Ord’s characterization of a true x-catastrophe as a single, defining event illustrates the notion of AI x-risk being about shockingly high-impact events. The decisive AI x-risk, in this context, is the expected uncertainty of the occurrence of such conclusive events as catalysts for x-catastrophic outcomes. [^5] Typically, the term ”risk” is quantitatively defined in terms of likelihood and impact. This quantitative approach involves assessing the probability of a specific event occurring and quantifying the potential consequences. However, in x-risk studies, researchers employ both qualitative and quantitative methods, each offering unique insights but also presenting specific limitations (Tonn and Stiefel, 2013; Beard et al., 2020). Quantitative methods for estimating AI x-risk require employing statistical models and subjective probability analysis. This approach is valuable for a seemingly structured way of estimating risks. However, it faces limitations due to the reliance on available data, which can be scarce or unreliable for unprecedented risks, and the difficulty of robustly encapsulating the multifaceted nature of existential threats in numerical terms. Qualitative methods involve non-numerical analyses such as scenario development, expert interviews, and ethical deliberations (Technology and Science Insights and Foresight, 2023). This approach is particularly useful in exploring the nuanced, complex, and often speculative aspects of AI risks, especially in areas where empirical data is scarce. However, it lacks numerical precision. This paper adopt a neutral stance regarding the quantitative or qualitative interpretation of AI x-risks. This neutrality allows me to draw on the broadest range of insights and perspectives in my analysis of AI x-catastrophes and their associated risks.
@@
Typically, existential and social risks are demarcated along the lines of locality — i.e., the scope of risk — and severity — i.e., hurting one or all including future generations (Bostrom, 2013; Amodei et al., 2016). While this demarcation is insightful, there is a notable gap in exploring the relationship between x-risks and the evolution of ethical concerns in a substantial manner. [^8] This disconnect might stem from two key reasons. First, discussions about decisive x-risks historically have revolved around reinforcement learning and agent-environment models, concentrating on existential threats from agency-based models. Ethical risks, in contrast, embrace a more expansive conceptual framework. Second, prominent conventional voices in ASI x-risks, such as Bostrom, Ord, and Yudkowsky, often subscribe to worldviews like rationalism, effective altruism, or longtermism. These prescriptive worldviews may be perceived as either problematic or tangential by some who are deeply engaged in the multifaceted discussion of ethical risks. The resultant divergence in normative viewpoints on guiding AI risk discourse and community priorities has led to a significant schism between these domains of risk (for a detailed analysis of such community divergences, see for example Ahmed et al. (2023)). A thorough exploration of these distinctions, however, falls outside the scope of this paper.
The social risks of AI systems have been extensively analyzed and studied, leading to the identification of six major risk categories: (1) lack of accountability and transparency about AIs and their owners, (2) representational and allocative unfairness and wrongful biases from AI predictions and automated decisions, (3) risk of the amplification of AI-generated misinformation, (4) threats to individual and group privacy, (5) malicious and dual use of AI, and (6) systemic risks impacting political, economic, and environmental stability. [^9] For lack of AI accountability and transparency, see Mitchell et al. (2019); Raji et al. (2020); Gebru et al. (2021); for representational and allocative unfairness and wrongful biases from AI predictions and automated decisions, see Buolamwini and Gebru (2018); Chouldechova and Roth (2020); Kleinberg et al. (2016); for the amplification risk of AI-generated misinformation, see Sharma et al. (2019); Quach (2020); Lin et al. (2021); for risks to individual and group privacy, see Dwork (2006); Tucker (2018); for risks of malicious and dual use, see Brundage et al. (2018); and for systemic risks impacting political economic, and environmental stability, see Korinek and Stiglitz (2018); Pashentsev (2021). For general taxonomies of the social risks of generative AI, see Weidinger et al. (2021); Bird et al. (2023); Barnett (2023). This paper posits that the gradual and cumulative progression of AI social risks delineates an alternative causal pathway to AI x-catastrophes:

### v3

@@ Abstract
The conventional discourse on existential risks (x-risks) from AI typically focuses on abrupt, dire events caused by advanced AI systems, particularly those that might achieve or surpass human-level intelligence. These events have severe consequences that either lead to human extinction or irreversibly cripple human civilization to a point beyond recovery. This discourse, however, often neglects the serious possibility of AI x-risks manifesting incrementally through a series of smaller yet interconnected disruptions, gradually crossing critical thresholds over time. This paper contrasts the conventional decisive AI x-risk hypothesis with an accumulative AI x-risk hypothesis. While the former envisions an overt AI takeover pathway, characterized by scenarios like uncontrollable superintelligence, the latter suggests a different causal pathway to existential catastrophes. This involves a gradual accumulation of critical AI-induced threats such as severe vulnerabilities and systemic erosion of economic and political structures. The accumulative hypothesis suggests a boiling frog scenario where incremental AI risks slowly converge, undermining societal resilience until a triggering event results in irreversible collapse. Through systems analysis, this paper examines the distinct assumptions differentiating these two hypotheses. It is then argued that the accumulative view can reconcile seemingly incompatible perspectives on AI risks. The implications of differentiating between these causal pathways — the decisive and the accumulative — for the governance of AI as well as long-term AI safety are discussed.
Forthcoming in Philosophical Studies
@@ 1 Introduction
Recent advances in machine learning have sparked intense debate about the existential risks (x-risks) associated with Artificial intelligence (AI) systems. [^1] For a review of recent narratives about AI and future, see Gilardi et al. (2024). Central to this debate is a concern about the potential pathways through which AI could cause existential catastrophes. In direct response to this concern, this paper explores: What are the distinct types of causal pathways through which AI systems could trigger existential catastrophes? Conventional discourse on AI existential catastrophes typically portrays them as sudden, decisive events, often triggered by artificial general or super intelligence (Bostrom, 2013) or extremely powerful (non-general) AI (Carlsmith, 2022).
Contrasting this conventional decisive viewpoint, this paper introduces accumulative AI x-risk hypothesis as an alternative lens. The accumulative hypothesis posits that AI x-risks do not exclusively materialize as decisive, high-magnitude global events initiated by extremely powerful AI such as artificial general or super intelligence. Instead, locally significant AI-driven disruptions can accumulate and interact over time, progressively weakening the resilience of critical societal systems, from democratic institutions and economic markets to social trust networks. When these systems become sufficiently fragile, a modest perturbation could trigger cascading failures that propagate through the interdependence of these systems. The failures amplify and reinforce each other through network effects and feedback loops, potentially leading to an irreversible civilizational collapse.
This paper develops an accumulative perspective on AI existential risk, by examining how multiple AI risks could compound and cascade over time to bring about an AI-generated existential catastrophe. Through applying systems analysis — an approach not typically used in AI x-risk scenario development — I defend the significance of accumulative AI x-risk, and consequently argue for a fundamental reconceptualization of AI x-risk governance, if such risks are to be governed effectively.
@@ 2 AI x-risk: preliminaries
2.1 Concepts of risk
At most basic, risk relates to some characterization of uncertainty about potential (adverse) outcomes (Dean, 1998; Hansson, 2010; Aven, 2012). According to the ISO 31000 standard (International Organization for Standardization, 2018), risk is defined as “the effect of uncertainty on objectives.” The Society for Risk Analysis (Aven et al., 2018) defines it as “uncertainty about and severity of the consequences of an activity,” while the U.S. Environmental Protection Agency (U.S. Environmental Protection Agency, 2024) defines risk as predicting “the probability, nature, and magnitude of the adverse effects that might occur.”
More concretely, risk has been analyzed through at least four distinct, though not mutually exclusive, interpretations.
First, risk has been used to mean an unwanted event that may occur (Carlsmith, 2022). For instance, “AI systems exhibiting power-seeking behavior pose a major existential risk to humanity.” Second, risk has been used to denote the cause of an unwanted event that may occur (Weidinger et al., 2022). For example, “Insufficient testing of generative AI models is a significant risk in AI deployment.” Third, risk has been used to mean the probability of an unwanted event that may occur (Lowrance, 1976). An example is “The risk that a large language model will generate incorrect information in response to user queries is approximately 5%.” Fourth, risk has been used to mean a statistical expectation value — the product of probability and consequence — of an unwanted event that may occur (United Nations International Strategy for Disaster Reduction, 2009). For instance, “The risk of LLM hallucination in a medical context can be calculated by multiplying the probability of giving incorrect medical advice (1%) by the average cost of medical liability claims (\$500,000), yielding an expected cost of \$5,000 per consultation.”
@@ 2.1 Concepts of risk
The particular focus of this paper is existential risks from AI. Existential risks (x-risks) refer to the potential for outcomes that would result in the extinction of humanity or an unrecoverable decline in humanity’s potential to thrive (Ord, 2020a). [^2] Some definitions, such as the one proposed by Bostrom (2013, p. 15), broaden the scope of existential threats to include not only human life but all sentient beings: “An existential risk is one that threatens the premature extinction of Earth-originating intelligent life or the permanent and drastic destruction of its potential for desirable future development.” Despite this broader definition, this paper deliberately employs the term ”humanity” as the focal subject of AI-induced existential catastrophes (Ord, 2020a). This choice is pragmatic and reflects the predominant narrative in AI x-risk scholarly literature and public discourse, which has traditionally concentrated on catastrophic impacts on humanity and human civilization. Additionally, this choice aligns with the stated missions of leading AI research companies like Google DeepMind and OpenAI in their pursuit of AGI or ASI development ”for the benefit of humanity.” My choice of terminology, however, does not diminish the moral significance of non-human sentient beings. Existential catastrophes are a class of potential events that may originate from natural causes, such as a supervolcanic eruption; anthropogenic sources, like nuclear conflict; or emerging threats, such as misaligned artificial superintelligence. [^3] For a historical examination of existential risks, see Torres (2023b). This paper concentrates on existential catastrophes induced by AI (AI x-catastrophes) and their associated risks.
In AI x-risk studies and discourse, researchers employ either of the qualitative and quantitative interpretations described above, each highlighting specific information about x-risks but also presenting specific limitations (Tonn and Stiefel, 2013; Beard et al., 2020). [^4] Quantitative methods for estimating AI x-risk require employing statistical models and subjective probability analysis. These methods are valuable for a seemingly structured way of estimating risks. However, they face limitations due to the reliance on available data, which can be scarce or unreliable for unprecedented risks, and the difficulty of robustly approximating the multifaceted nature of existential threats from AI in numerical terms. Qualitative methods involve non-numerical analyses such as scenario development, expert interviews, and ethical deliberations (Technology and Science Insights and Foresight, 2023). The qualitative methods are particularly useful in exploring the nuanced, complex, and often speculative aspects of AI risks, especially in areas where empirical data is scarce. This paper aims to maintain neutrality between these different interpretations of x-risk, drawing on a broad range of perspectives to analyze AI catastrophes and their associated risks. While I employ the causal interpretation of AI existential risks for illustrative purposes, alternative interpretations could remain equally valid. The probabilistic and statistical interpretations of x-risk warrant their own detailed investigation elsewhere.
@@
The conventional discourse on AI x-catastrophes portrays them as decisive, large-scale events caused by highly advanced AI systems, often referred to as artificial general intelligence (AGI) or artificial superintelligence (ASI). [^5] In AI lexicon, AGI represents capabilities comparable to human intelligence, while ASI denotes capabilities surpassing human intelligence. Traditional arguments typically link x-risks primarily to AGI or ASI (e.g., Bostrom (2014); Ord (2020a, b)), though some suggest that AI x-risks could also emerge from AI systems with extreme but narrow capabilities (Carlsmith (2022)). This paper’s focus on AGI/ASI aligns with influential voices in x-risk studies such as major AGI research companies. However, my core argument about the nature of x-risks holds whether we consider AGI/ASI or narrower AI systems with extreme, uncontrollable capabilities in specific domains. Thus, this terminological choice does not affect my characterization of conventional views on AI x-risk. The idea that advanced machines can pose significant risks is not novel and has historical antecedents. [^6] For a comprehensive historical review, see Torres (2023a). Samuel Butler (1863, p.185), a novelist and literary critic, alluded to the possibility of machines dominating humanity. This concern was later picked up by the renowned mathematician Alan Turing (1950), who warned of intelligent machines eventually taking control. Norbert Wiener (1960), a founder of the field of cybernetics, cautioned against entrusting machines with purposes misaligned with human intentions or desires. Similarly, the mathematician Irving J. Good (1966) expressed concerns about the creation of “ultraintelligent machines.”
@@
More recently, philosophers like Nick Bostrom (2002, p.7) drew systematic attention to the existential threats posed by ASI: “When we create the first superintelligent entity, we might make a mistake and give it goals that lead it to annihilate humankind, assuming its enormous intellectual advantage gives it the power to do so.” [^7] While Bostrom’s work has systematized the discussion of x-risks from ASI, the original post-2000 discussions trace back to Eliezer Yudkowsky’s views concerning AGI, ASI, and their associated x-risks. Ben Goertzel (2015) explores the early conceptual evolution of these topics, tracing their roots to Yudkowsky’s initial informal explorations. In this paper, I primarily reference Nick Bostrom as a representative figure who brought systematic and philosophical depth to concepts that Yudkowsky and others initially introduced and examined in speculative media and blog posts. Computer scientists such as Stuart Russell (2019) and physicists like Max Tegmark (2018) echoed similar concerns, stressing the x-risks of ASI beyond human control. Such views about ASI x-risks frequently hinge on two key theses: orthogonality and instrumental convergence (Bostrom, 2012, 2014). [^8] For a critical discussion of the orthogonality and instrumental convergence theses, see Müller and Cannon (2022).
The orthogonality thesis posits that an advanced AI system’s intelligence level and its final goals are orthogonal. [^9] See Bostrom (2012, p.73) and Bostrom (2014, p.107). This implies that an AI system could have any combination of final goals, beneficial or harmful, and intelligent capabilities. The instrumental convergence thesis holds that diverse final goals often have similar instrumental sub-goals — like self-preservation or resource acquisition — as these sub-goals are useful for achieving almost any final objective. An ASI might therefore pursue actions harmful to humanity not out of malice, but as instrumental steps toward achieving its programmed objectives. Two hypothetical scenarios illustrate how conventional models of ASI development could lead to x-catastrophic outcomes.
In a thought experiment popularized by Bostrom (2003), an ASI is given the seemingly innocuous and simple goal of maximizing paperclip production. Even with this simple objective, the ASI could pursue instrumental sub-goals: it might eliminate humans to prevent deactivation (self-preservation) or convert their bodies into paperclips (resource acquisition). Consequently, the ASI’s optimal future could become one abundant with paperclips but devoid of humans — not because intelligence necessitates this goal, but because instrumental sub-goals are (putatively) rational steps toward satisfying its given optimization objective. That is, the paperclip maximizer illustrates how an AI with an apparently harmless goal could pose an x-risk through the rational pursuit of instrumental sub-goals like resource acquisition and self-preservation. [^10] The paperclip maximizer is characterized by some as a “Squiggle minimizer”: a highly intelligent optimizer pursuing goals alien to human values could inadvertently destroy humanity by consuming vital resources in pursuit of its objectives. See LessWrong (2024).
In a structurally similar thought experiment, Russell and Norvig (2010, p.1039) describe a scenario attributed to Marvin Minsky: an advanced AI tasked with proving the Riemann hypothesis might appropriate Earth’s resources to build supercomputers, endangering humanity in pursuit of a mathematical proof. Like the paperclip maximizer, this example is supposed to illustrate how an advanced AI pursuing a benign goal could pose existential risks to humanity. Both scenarios demonstrate a key pattern: an advanced AI system optimizing for a specific goal could rationally pursue sub-goals which are catastrophic to humanity, not from malice but as instrumental steps toward its optimization goal.
2.2 Decisive AI x-risk
@@
The conventional view, sketched above, frames AI x-catastrophes arising through decisive actions of ASI. Toby Ord (2020a, p. 20), a prominent x-risk scholar, explicitly endorses the decisive character of x-catastrophes: “I take on the usual sense of catastrophe as a single decisive event rather than any combination of events that is bad in sum. A true existential catastrophe must by its very nature be the decisive moment of human history the point where we failed.” This conventional framing suggests that AI x-risks manifest as sudden, cataclysmic events that either eradicate humanity or irreversibly curtail its potential. I articulate this perspective in terms of the decisive ASI x-risk hypothesis:
@@
Both Bostrom’s portrayal of ASI pursuing destructive goals and Ord’s characterization of existential catastrophes as singular, defining events emphasize the conventional view: AI x-risks manifest as sudden, decisive moments of overwhelming impact. The decisive AI x-risk, according to this framing, is the expected uncertainty of the occurrence of such conclusive events as catalysts for x-catastrophic outcomes.
@@ 2.3 AI risks: existential versus social
Typically, existential and social risks are demarcated along the lines of locality — i.e., the scope of risk — and severity — i.e., whether impacts are recoverable and limited, or irreversible and catastrophic for (human) civilization (Bostrom, 2013; Amodei et al., 2016). While this demarcation is pragmatically insightful, there is a notable gap in exploring the relationship between x-risks and the evolution of ethical concerns in a substantial manner. [^13] This disconnect might stem from two key reasons. First, discussions about decisive x-risks historically have revolved around reinforcement learning and agent-environment models, concentrating on existential threats from agency-based models. Ethical risks, in contrast, embrace a more expansive approach, covering non-reinforcement approaches to AI development. Second, prominent conventional voices in ASI x-risks, such as Bostrom, Ord, and Yudkowsky, often subscribe to worldviews like rationalism, effective altruism, or longtermism. These prescriptive worldviews may be perceived as either problematic or tangential by some who are deeply engaged in the multifaceted discussion of ethical risks. The resultant divergence in normative viewpoints on guiding AI risk discourse and community priorities has led to a significant schism between these domains of risk (for a detailed analysis of such community divergences, see for example Ahmed et al. (2023)). A thorough exploration of these distinctions, however, falls outside the scope of this paper.
The social risks of AI systems have been analyzed across different domains — from language models and their multimodal variants (Bender et al., 2021; Weidinger et al., 2022; Bird et al., 2023) to recommender systems (Milano et al., 2020; Deldjoo et al., 2024), to name just a few examples. While a full exposition of these risks is beyond the scope of this paper and has been extensively discussed in the above references, here I provide a brief categorization.
Manipulation and deception risks include AI systems that cause harm by manipulating human behavior through targeted or unwanted persuasion (Kasirzadeh and Evans, 2023; Carroll et al., 2024). Misinformation and disinformation risks arise from AI systems generating and amplifying false content at scale, enabling the spread of propaganda and undermining public trust and discourse (Sharma et al., 2019; Quach, 2020; Lin et al., 2021; Kay et al., 2024). Malicious use risks include the weaponization of AI systems for cyber attacks, the deployment of AI-enabled drones and other physical systems for attacks, as well as the automation of social engineering attacks (Brundage et al., 2018). Insecurity and information threat risks arise when AI systems reveal sensitive personal data or lead to an unintended disclosure of protected information (Carlini et al., 2021). Discrimination and hate speech risks manifest through biased AI systems perpetuating systemic inequalities or generating targeted harmful content (Buolamwini and Gebru, 2018; Obermeyer et al., 2019). Surveillance, rights infringement, and erosion of trust risks originate from AI-powered mass surveillance systems and persistent monitoring that could lead to loss of privacy (Dwork, 2006; Tucker, 2018) and loss of trust in ruling institutions (Nowotny, 2021). Environmental and socioeconomic risks result from massive energy consumption of AI training or displacement of workers through automation (Korinek and Stiglitz, 2018; Pashentsev, 2021).
2.4 Accumulative AI x-risk
As an alternative to the decisive AI x-risk hypothesis, the gradual and cumulative progression of critically significant social risks characterizes a different type of causal pathway to AI x-catastrophes: